# Gauss-Newton accelerated MPPI Control

### v1

@@ 4.1 Jacobian reconstruction
$\mathrm{cost}_{\mathrm{GPU}}\left(\frac{\partial\hat{R}}{\partial U}(U_{k})\right)\approx\mathrm{cost}(R(U_{k}))$ (12)
follows under the hard assumption of ideal parallelization. We can approximate the Jacobian $J(U_{k}):=\frac{\partial\hat{R}}{\partial U}(U_{k})$ of the smoothed inner black-box function $R$ of (4). Note that in the case of a differentiable inner function $R$, finite differences (Curtis.1974) is another approach to approximate the Jacobian.
@@ 4.3 Gauss-Newton accelerated MPPI algorithm
where $\mathcal{A}=\{\gamma^{-k}|k=0,\ldots,M-1\}$ is a set containing $M$ applicable step lengths with $\gamma\in(0,1)$. The computation of the optimal step length can be executed in parallel, leading to a computational cost
@@
$\mathrm{cost}_{\mathrm{GPU}}\left(\mathrm{Algo.~1}\right)\approx 2K\cdot\mathrm{cost}(R(U_{k})).$ (17)
Note that the computational cost estimates are derived under the assumption of ideal parallelization and that the computational cost of the simulator Eq. (1) dominates that of all other algorithmic components. To satisfy the first assumption, the number of samples $M$ must be chosen to be smaller than about a few thousand.
@@ 5 Numerical Experiments
Method A: GGN with finite difference approximation of the Jacobian of $R$ in forward mode with small perturbations $\epsilon\approx 1.49\times 10^{-8}$ (Curtis.1974, Eq. 2.2).
@@
Method D: Gauss-Newton accelerated MPPI, as presented in Algorithm 1.
For Method C and Method D, $M=2000$ samples are used, and antithetic sampling is employed to improve performance, cf. Sec. 4.2 in Glasserman.2003. For reproducibility, the full implementation is provided in the repository: tba -- after publication.
@@ 5.1 Results
The optimization is stopped if the step length and the (smoothed) gradient of the objective fall below a small threshold. For each problem, the method that achieved the optimal solution using the fewest iterations is highlighted. The iteration count is a suitable indicator of computational effort, since each method requires two sequential evaluations of the inner function.
@@
The results show that the GGN method with finite differences (Method A) achieves the fastest convergence in well-posed problem settings such as the Rosenbrock function and linear or nonlinear MPC with smooth objective functions. However, it converges only to a suboptimal solution for the Rastrigin function and is not admissible for optimizing nonsmooth problems with vanishing gradients. The gradient descent method with finite differences (Method B) generally requires a large number of iterations to converge (see, however, Problem II). It fails to solve the Rosenbrock function within 1000 iterations.
Both the deterministic MPPI method (Method C) and the Gauss-Newton accelerated MPPI method (Method D) converge across all tested problem classes. In particular, the Gauss-Newton accelerated MPPI method consistently achieves faster convergence (up to $\times$9 fewer iterations), especially in high-dimensional settings (cf. Problems IV, V.i, and V.ii). Note that the deterministic MPPI method requires 45 iterations for the linear optimal control example (Problem IV), even though this problem is a quadratic program that could be solved in a single Newton step. Remember the limitation of the presented Gauss-Newton accelerated MPPI approach given by the restriction to the convex-over-unknown problem structure (4). This structure is not necessary for the generalized deterministic MPPI approach. However, as previously discussed, this structure is widespread in optimal control applications.

### v2

@@ 4.1 Jacobian reconstruction
$\mathrm{cost}_{\mathrm{GPU}}\left(\frac{\partial\hat{R}}{\partial U}(U_{k})\right)\approx\bigg\lceil\frac{M}{M_{\mathrm{GPU}}}\bigg\rceil\cdot\mathrm{cost}(R(U_{k}))$ (12)
follows under the hard assumption of ideal parallelization of $M_{\mathrm{GPU}}\in\mathbb{N}$ processing units. We can approximate the Jacobian $J(U_{k}):=\frac{\partial\hat{R}}{\partial U}(U_{k})$ of the smoothed inner black-box function $R$ of (4). Note that in the case of a differentiable inner function $R$, finite differences (Curtis.1974) is another approach to approximate the Jacobian.
@@ 4.3 Gauss-Newton accelerated MPPI algorithm
where $\mathcal{A}=\{\gamma^{k}|k=0,\ldots,M-1\}$ is a set containing $M<M_{\mathrm{GPU}}$ applicable step lengths with $\gamma\in(0,1)$. The computation of the optimal step length can be executed in parallel, leading to a computational cost
@@
$\mathrm{cost}_{\mathrm{GPU}}\left(\mathrm{Algo.~1}\right)\approx 2K\cdot\bigg\lceil\frac{M}{M_{\mathrm{GPU}}}\bigg\rceil\cdot\mathrm{cost}(R(U_{k})).$ (17)
Note that the computational cost estimates are derived under the assumption that the computational cost of the simulator Eq. (1) dominates that of all other algorithmic components.
@@ 5 Numerical Experiments
Method A: GGN with central finite difference approximation of the Jacobian of $R$ in mode with small perturbations $\epsilon\approx 1.49\times 10^{-8}$ (Berahas.2022, Eq. 1.2).
@@
Method D: Gauss-Newton accelerated MPPI (cf. Algo. 1) with random sampling of Eq. (10).
Method E: Gauss-Newton accelerated MPPI (cf. Algo. 1) with deterministic Dirac-mixture approximation of Eq. (10). This technique is based on the sigma-point approach utilized in the Unscented Kalman Filter (UKF) (Wan2000).
For Methods C and D, $M=2000$ samples with $\lceil M/M_{\mathrm{GPU}}\rceil=1$ are used, and antithetic sampling is employed to improve performance, cf. Sec. 4.2 in Glasserman.2003. For reproducibility, the full implementation is provided in the repository: tba -- after publication.
@@ 5.1 Results
The optimization is stopped if the step length and the (smoothed) gradient of the objective fall below a small threshold. The iteration count is a suitable indicator of computational effort, since each method requires two sequential evaluations of the inner function.
@@
The results show that the GGN method with finite differences (Method A) achieves fast convergence in well-posed problem settings such as the Rosenbrock function and linear or nonlinear MPC with smooth objective functions. However, it converges only to a suboptimal solution for the Rastrigin function and is not admissible for optimizing nonsmooth problems with vanishing gradients. The gradient descent method with finite differences (Method B) generally requires a large number of iterations to converge (see, however, Problem II). It fails to solve the Rosenbrock function within 1000 iterations.
In contrast, the deterministic MPPI method (Method C) and the Gauss-Newton accelerated MPPI methods (Variant D and E) converge across all tested problem classes. In particular, Method E consistently achieves faster convergence (up to $\times$21 fewer iterations), especially in high-dimensional settings (cf. Problems IV, V.i, and V.ii) compared to deterministic MPPI (Method C). Note that the deterministic MPPI method requires 45 iterations for the linear optimal control example (Problem IV), even though this problem is a quadratic program that could be solved in a single Newton step. Remember the limitation of the presented Gauss-Newton accelerated MPPI approach given by the restriction to the convex-over-unknown problem structure (4). However, as previously discussed, this structure is widespread in optimal control applications.